OpenStudy (anonymous):

y=x^2+14x+50 in vertex form. Help explain this to me? both people who have answered have different answers which confuses me.

OpenStudy (anonymous):

i think it might be(7,1). don't know for sure

OpenStudy (anonymous):

was i right

OpenStudy (anonymous):

Still solving it, but i think you are

OpenStudy (anonymous):

the correct answer is -7 ,-1

OpenStudy (anonymous):

It's 7,1. The equation is y=x^2+14+50. may i see your work?

OpenStudy (anonymous):

@AJ1400

OpenStudy (anonymous):

Write the equation as follows \[ y= x^2+14x+50= x^2 + 14 x + 49 +1=\\ y=(x+7)^2 +1 \] So the vertex is (-7,1) See also the wolframalpha plot